An aliquot of 10 mL from a 12-mol/L aqueous silver solution was transferred into a 100-mL volumetric flask and diluted with water up to the mark. A 10-mL aliquot of the diluted silver solution was transferred into another 100-mL volumetric flask and diluted with water up to the mark. What is the concentration of the final dilution?


1
Expert's answer
2022-02-23T05:46:04-0500

C1 = 12 mol/L

V1 = 10 mL

C2 = unknown

V2 = 100 mL

V2" = 10 mL

Cf = unknown

Vf = 100 mL


Solution:


Therefore, after the first dilution, the concentration of the solution (С2) will be:

С2 = (С1 × V1) / V2 = (12 mol/L × 10 mL) / (100 mL) = 1.2 mol /L


Calculate the concentration of the solution after the final dilution:

С2 × V2" = Cf × Vf

Cf = (С2 × V2") / Vf = (1.2 mol /L × 10 mL) / (100 mL) = 0.12 mol /L

Cf = 0.12 mol /L


Answer: The concentration of the final dilution is 0.12 mol/L